QN.The XYZ Company Ltd is facing an option to develop 2 models of the same product: Model A and Model B. Both models are expected to be hot sellers for the next two years. The capital expenditure involved in the development of model A is Tshs 2,250 million while Model B costs Tshs 2500 Million. Because of familiarity of the existing users, Model A (that is an improved version of the existing one) offers a certain cash flow of Tshs 1,400 million and 1,800 million for the next two years.
In respect of Model B, however, XYZ Ltd is not very sure of the cash inflows. It expects a cash inflow of Tshs 1,500 million in the first year with 70% possibility. Under exceptional circumstances the product may do better to yield a cash inflow of Tshs 1,800 million. The cash inflows in the second year are dependent upon what happens in the first year. If the firm faces a normal cash inflow of Tshs 1,500 million in the first year the chances are that it would repeat with probability of 30%, improve to Tshs 1,800 million with 40% chance and do exceptionally well with cash inflow of Tshs 2,200 million having 30% probability. If the firm faces better prospects in the first year, the chances are that it would have second year cash inflows of Tshs 2,200 million, 2,300 million and Tshs 2,500 million with probabilities of 50%, 30% and 20% respectively.
Required:
Examine both models and recommend to XYZ Ltd which of the two models they must go ahead with. Assume a cost of capital of 12%.
Answer #1

All financials below are in Tshs million.

NPV of model A = -C0 + PV of all future cash flows = -2,250 + 1,400 / (1 + 12%) + 1,800 / (1 + 12%)2 =  434.95

Model B

Expected cash flows in year 1 = 1,500 x 70% + 1,800 x 30% = 1,590

Expected cash flows in year 2 = 70% x (30% x 1,500 + 40% x 1,800 + 30% x 2,200) + 30% x (50% x 2,200 + 30% x 2,300 + 2,400 x 20%) = 1,962

Hence, NPV of model B = -C0 + PV of all expected future cash flows = -2,500 + 1,590 / (1 + 12%) + 1,962 / (1 + 12%)2 =  483.74

NPV of model B > NPV of model A

Hence, they must go ahead with model B